40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 18 00:58:40 2011, jamie, Update, SUS, optics kicked and left free swinging 
    Reply  Thu Aug 18 12:22:37 2011, jamie, Update, SUS, suspension update 6x
Message ID: 5260     Entry time: Thu Aug 18 00:58:40 2011     Reply to this: 5263
Author: jamie 
Type: Update 
Category: SUS 
Subject: optics kicked and left free swinging 

ALL optics (including MC) were kicked and left free swinging at:

997689421

The "opticshutdown" script was also run, which should turn the watchdogs back on in 5 hours (at 6am).

 

ELOG V3.1.3-